OpenStudy (anonymous):

The formula for a geometric sequence is an = 1000(0.5)^n-1. What is the 3rd term of the sequence?

OpenStudy (anonymous):

what do you mean by an? - the nth term?

OpenStudy (anonymous):

\[a _{n}\]

OpenStudy (anonymous):

Assuming this answer is correct I found this online, it's the same question. (last question asked) http://answers.yahoo.com/question/index?qid=20101122113057AAlV9Ld

OpenStudy (anonymous):

Thanks PsychoTink! :)

OpenStudy (anonymous):

yes if we put x=3 in your formula 3rd term is 1000*0.5^2 = 1000 * 0.25 = 250

OpenStudy (anonymous):

Thanks jimmyrep! :)

OpenStudy (anonymous):

Answer is 250

OpenStudy (anonymous):

Oh, That makes sense. Cool

OpenStudy (anonymous):

Thank you SweetRascal! :)

OpenStudy (anonymous):

You are welcome :)

OpenStudy (anonymous):

yw